-If the value of chi-square is large and cannot be accounted for based on sampling error, the researcher should:

fail to reject the null hypothesis

reject the null hypothesis

determine the degrees of freedom in the study

observe the chi-square in the common zone.

-If %Expected = 19 and N = 33, what is the value of fExpected?

6.08

0.63

6.27

5.7

-A psychology professor is studying whether a significant difference exists in the majors among students in her general psychology course. She splits majors into humanities, sciences, and business. If there are 75 students in her class, and the expected value of humanities majors is 16.7 and the expected value of sciences is 47.8, what is the expected frequency of business majors?

12.53

35.85

7.88

10.50

-A researcher is conducting a chi-square goodness-of-fit test on a variable that has three categories. The observed frequencies are 25 for Cell A, 54 for Cell B, and 37 for Cell C. The expected frequencies are 31.56, 47.78, and 36.66 for Cells A, B, and C, respectively. What is the value of chi-square?

1.37

0.81

2.18

2.53
Answer #1

1) fExpected =np =33*0.19 =6.27

2) expected frequency of business majors =75-16.7-47.8 =10.5

3)

chi-square =Σ(Oi-Ei)^2/Ei =(25-31.56)^2/31.56+(54-47.78)^2/47.78+(37-36.66)^2/36.66 =2.18